OpenStudy (anonymous):

The measure of an interior angle of a regular polygon is 147 3/11. Find the number of sides in the polygon

OpenStudy (lgbasallote):

the formula for interior angle is \[\large \frac{180(n-2)}{n}\] where n is the number of sides so you say the interior angle is 147 3/11 \[\frac{180(n-2)}{n} = 147 \frac{3}{11}\] \[\frac{180(n-2)}{n} = \frac{1620}{11}\] \[1980(n-2) = 1620n\] \[1980n - 3960 = 1620n\] \[1980n - 1620n = 3690\] solve for n

OpenStudy (anonymous):

10.25?

OpenStudy (lgbasallote):

seems like it

OpenStudy (lgbasallote):

you're right

OpenStudy (anonymous):

oh thank u so do i have to estimate to the nearest tenth so it would be just 10?

OpenStudy (lgbasallote):

depends on your teacher's preference

OpenStudy (lgbasallote):

normally that's fine as is